UT 20130915 Technical Observing - AGw2 Calibration

Observer: JHill, DMiller (Tucson)
Telescope Operator: SAllanson (LBTO)
Software Support: DCox (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: (LBTO)
Instrument Support: JPower (LBTO)

Summary

Lost to weather: Clouds, humidity and nearby thunderstorms kept us closed at the start of the night, and the humidity never went away. It eventually rained around 09:30 UT.

John spent much of the night developing a Python script to take symbiotic LBC-Red data in PseudoBinocular mode. (Issue 4513)

Olga spent much of the night diagnosing a problem with MODS1 red collimator. (Issue 4801)

Preparations

Details

-- JohnHill - 15 Sep 2013
Topic revision: r4 - 16 Sep 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback